OpenStudy (anonymous):

Adult IQ scores have a bell-shaped distribution with a mean of 100 and a standard deviation of 15. Use the Empirical Rule to find the percentage of adults with scores between 85 and 115.

OpenStudy (anonymous):

Z_1 = ( 85 - 100 ) / 15 = ... Z_2 = ( 115 - 100 ) / 15 = ... Look up the table for probabilty

OpenStudy (anonymous):

is this the answer 68%

OpenStudy (anonymous):

Yupe..eee! P ( -1 < z < 1 ) = 68%

OpenStudy (anonymous):

i need help with another

OpenStudy (anonymous):

Just open new post as you've done so far!